OpenStudy (anonymous):

(-4-1)^2. Do I foil this or add -4 and -1 and square that?

OpenStudy (campbell_st):

well your problem has lots of solutions... here is 1 \[(-4 -1)^2 = (-4 -1)\times (-4 -1) \] but the easiest is the add the numbers -4 -1 = -5 now square the answer.

OpenStudy (anonymous):

Alright. Thanks!